40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed Jul 15 23:17:14 2015, Alessandra, Misc, Seismometer, Rhomboid motion Sensors.JPGfoto_.JPGfoto_(1).JPG
    Reply  Thu Jul 16 10:21:39 2015, Koji, Misc, Seismometer, Optical lever installation (Re: Rhomboid motion) 
Message ID: 1979     Entry time: Thu Jul 16 10:21:39 2015     In reply to: 1978
Author: Koji 
Type: Misc 
Category: Seismometer 
Subject: Optical lever installation (Re: Rhomboid motion) 

Some supplimental info on Alessandra's entry

0. Suspension stopper: In order to make the work on the suspension easier, I made simple suspension stoppers.

1. Oplev layout: We decided to put all of the input-output optics on a single triangle platform so that the optical lever angle become small. However, this was a challenge as the x-z stage for the QPD calibration is bulky... Once the calibration is done, the stage can be replaced with a pole.

2. The laser: The bare outputs of the laser diodes are too dirty for the optical lever measurement. We decided to use a fiber pigtail laser as the mode quality is excellent.We needed to setup a collimation optics. I brought a bunch of fiber coupling optics including an aspherical collimation lens with f=10mm. It was adjusted to have the focus at ~1m from the collimator. The beam radius looked like ~0.5mm. Allessandra will check this this afternoon.

3. The reflection mirror: Since the beam height on the laser platform determines the optical height at the rhomboid that is higher than the middle plate of th rhomboid, we needed to use an optical mounts on the middle plate. In fact we installed two (almost) identical mirror assemblies for counter balancing. This actually made the rhomboid unstable as the center of mass became higher than the clamping point. We decided to add a balast mass at the bottom of the rhomboid. This also work as a tilt adjuster. The balast mass was attached to the plate using a double sided tape. As a result, we recovered a stable condition (positive spring constant) of the pendulum, however, the tilting (pitching) frequency is now not precisely tuned. Maybe this is what Alessandra can work on after the QPD calibration?

ELOG V3.1.3-